A gaseous mixture of S2 and co2 gas contain 66% mass of co2. the vapour density of the mixture :A. 6.1B. 5.4C. 2.7 D. 10.8

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Akash
 
H2 should be in place of S2. Otherwise this question is wrong according to the options.
 
Let total moles of mixture be 100g
Moles of CO2 = 66g
Then hydrogen will be = 34g
Number of moles of CO2=66/44=1.5
Numbsr of moles if H2=34/2=17
Total moles =1.5+17=18.5
Formula to be used:
Mavg= Total mass/ total moles
= 100/18.5=5.40
VAPOUR DENSITY= MAvg/2
=5.40/2=2.70
So, VAPOUR density of mixture is 2.7
 
 
 
Regards
Arun (askIITians forum expert)